# Contrastive-KAN: A Semi-Supervised Intrusion Detection Framework for Cybersecurity with scarce Labeled Data

### v1

@@ 3.1 UNSW-NB15 Dataset
The UNSW-NB15 dataset created by researchers at the University of New South Wales (UNSW) represents modern network traffic scenarios and includes a wide range of low-foot print attacks. A synthetic testbed was configured using the IXIA PerfectStorm tool, that generated realistic network traffic, and a variety of attack scenarios. This dataset contains network traffic collected from multiple sources, including safe communications, malicious attacks, and routine operations such as software updates. Raw network packets were captured using the tcpdump tool and stored in PCAP file format.
A total number of 49 network-based features were extracted in five categories: basic features, content features, time features, connection features, and additional features, such as protocol type, duration, packet/byte counts, payload analysis, jitter, inter-packet arrival times, and historical behavior. This dataset includes multiple modern attack families: Fuzzers, Analysis, Backdoors, DoS, Exploits, Reconnaissance, Shellcode, Generic, and Worms, in addition to Normal traffic. In this work, we utilize only a subset of categories: Normal, Fuzzers, DoS, Exploits, Generic and Reconnaissance. The total number of records is about 2.5 million, consisting of 2,218,761 normal and 321,283 attack records [36].
The UNSW-NB15 dataset was collected to address the limitations of previous datasets (e.g. KDD99/NSL-KDD datasets), such as lack of modern low-footprint attacks and outdated normal traffic patterns. Skewness and Kurtosis tests on this dataset confirm non-linearity, and non-normal distributions and feature correlations that challenge traditional ML methods [38]. Older datasets like KDD99 and NSL-KDD suffered from redundant records. The UNSW-NB15 dataset addresses this issue [39]-[40]. The various classes and the corresponding number of instances for each condition are provided in Table 1.
It is worth noting that the original dataset contains 49 columns, including several categorical features: proto, service and state. We apply one-hot encoding to these features, this will result in a total of 204 columns after feature encoding. In this work, 558,954 samples are used for pre-training phase and 55,895 samples for fine-tuning stage, which corresponds to 22.01% and 2.20% of total number of samples, respectively. The fine-tuning set is sampled from the pre-training set.
@@ 3.2 BoT-IoT Dataset
The BoT-IoT dataset [37], publicly released in 2018 by the Network Security Lab at the University of New Brunswick, Canada, was collected using a smart home testbed that included various IoT devices like routers, cameras, and smart appliances. The setup was designed to mimic real-world IoT environments and includes multiple different types of cyber-attacks. Network traffic was recorded over several weeks using a passive monitoring tool called CICFlowMeter, which captured data at the network gateway.
This dataset contains network features, including protocol type, source/destination IP addresses, and port numbers, and each data point is labeled. It serves as a benchmark for evaluating intrusion detection systems, especially those focused on IoT security, and includes common IoT-related attacks. The categories present in this dataset: DDoS, DoS, Normal, Theft and Reconnaissance.
In practice, when preparing the data for training, some columns which are attack and subcategory are removed, however category is retained to serve as the target label. The proto column, being a categorical variable, is transformed using one-hot encoding, similar to preprocessing steps applied to the UNSW-NB15 dataset. After encoding, the dataset ends up with 20 features for training. For this dataset, we use 469,570 samples which are approximately 12.8% of the data for the pre-processing step, while for the fine-tuning phase 46,957 samples are employed which account for 1.28% of the total 3,668,522 samples. The fine-tuning subset is sampled from the pre-training subset of data, and we only use labels for the fine-tuning subset. Both numbers are significantly lower than the volume of data typically required by other algorithms. Table 2 presents the classes in BoT-IoT dataset along with the number of instances for each condition.
@@ 3.3 Gas Pipeline Dataset
The gas pipeline dataset is developed to study the intrusion detection in critical infrastructure systems that relies on SCADA technology. Its primary purpose is to simulate cyber-attacks targeting SCADA systems, serving as a standardized and common benchmark for evaluating different intrusion detection methods. This dataset is based on a small-scale gas pipeline testbed, where MODBUS traffic was recorded using a network data logger. The testbed is consists of an airtight pipeline connected to a compressor, a pressure sensor and a solenoid-controlled relief valve. The system regulates air pressure within the pipeline through a proportional-integral-derivative (PID) control scheme. In this research, we utilize the 2014 version of the dataset [35].
The dataset is composed of 27 columns, the result column, denotes the condition of the system. Therefore, the dataset actually has 26 actual features related to network traffic, process control and process measurements. Among these, seven features contain only a single constant value and do not contribute any useful information for the algorithm. These features are, rate, comm_write_fun, gain, reset, deadband, cycletime and crc_rate. After dropping these features, we are left with 19 features, that are going to be used for the training of the network. Regarding the labels, there are eight distinct conditions in the dataset, which are normal (Normal), Naïve Malicious Response Injection (NMRI), Complex Malicious Response Injection (CMRI), Malicious State Command Injection (MSCI), Malicious Parameter Command Injection (MPCI), Malicious Function Code Injection (MDCI), Denial Of Service (DoS) and Reconnaisance (Recon). These are denoted by labels, 0 to 7 respectively [35]. The classes along with their respective instance counts are shown in Table 3.
A total of 77,615 samples, that represents 80% of the entire dataset, are used for the pre-training phase. From this set, 7,761 samples (8% of the total data) are selected for fine-tuning. The pre-training sets require no labels, and labels are only used in the fine-tuning stage, making this approach suitable for scenarios with abundant unlabeled data.
@@ 4 Methodology
In this section, we present the methodology proposed in this study. In section subsection 4.1 begins with an overview of the Kolmogorov–Arnold representation theorem and KAN, subsection 4.2 outlines the pre-processing step associated with the method, and finally, in subsection 4.3, contrastive learning is thoroughly investigated.
@@ 4.1 KAN: Kolmogorov-Arnold Network
KAN, first introduced in [11], is an alternative to traditional fully-connected network. It exhibits strong capabilities in modeling non-linear functions and outperforms conventional MLP layers. KAN has demonstrated significant potential in a variety of domains, including, anomaly detection [41], human activity recognition [42], solving differential equations [43], time series analysis [44]. KAN has shown notable results, such as, advancement in accuracy, interpretability, and computational efficiency [11]. Before proceeding to the methodology proposed, we examine the current applications of KANs in the field of anomaly detection. [45] employs KAN for CPS attack detection, using a minimal and fast approach, which only utilizes standard-scaling, this method achieves high performance and low inference time.
In [46], the authors propose faster-KAN with the help of Gaussian radial basis functions (GRBF) to detect anomalies in IoT networks. This research is suitable for low-end devices and systems with limited computational resources. [41] introduces CCTAK, an unsupervised anomaly detection framework for IIoT networks, with a combination of CNN and KAN. The enhanced interpretability of this method is a key advantage. [47] uses KAN for credit card fraud detection. In this work, KAN outperforms the conventional MLP networks. In [48], the detection of botnet attacks is investigated. It is shown that the original KAN outperforms the traditional ML and DL methods, while also providing improved interpretability. [49] utilizes KAN for motor condition monitoring. KAN improves the real-time diagnosis of faults in the motors, achieving both better performance and interpretability compared to MLP networks. KAN is also adopted in healthcare domain. In [50], the authors performed the ECG-based cardiac abnormality detection, showcasing the effectiveness and generalizability of KAN for wearable medical devices.
KAN is grounded in the Kolmogorov–Arnold representation theorem, which initially served to solve differential equations. However, in this study, KAN is applied for classification tasks. The theorem guarantees that any multivariate continuous function $f(X)$, where $X=(X_{1},\dots,X_{T})$, can be decomposed into a sum of nested univariate functions. Meaning, for the smooth differentiable function $f:[0,1]^{T}\rightarrow\mathbb{R}^{d_{out}}$, there exist continuous univariate functions $\Phi_{q}$ and $\phi_{q,p}$, that satisfy the following equation:
@@
where $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\Phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. In this formulation, each $\phi_{q,p}$ and $\Phi_{q}$ is a univariate function, and the nested sum of them enables the approximation of any continuous multivariate function. The Kolmogorov-Arnold representation theorem ensures $(2T+1)\times T$ uni-variate functions are enough to approximate any given $T$-variate function. Compared to Taylor or Fourier series approximations, that require an infinite number of terms to represent a multivariate function with high precision, KANs offer a significant advantage by approximating a function using a finite composition of learnable uni-variate functions. KANs leverage this theoretical foundation by replacing typical pre-defined activation functions in neural networks with learned uni-variate functions applied to each input dimension individually. Unlike in a traditional MLP layer, where activation functions are applied to the outputs of nodes, KAN places activation functions on the edges of the network, i.e., between connected nodes. Figure 1 visualizes a KAN with an input size of $T$ and an output size of $C$. From this point on, to model the approximated function $f(\cdot)$, we use the notation $KAN(\cdot)$. And $X^{i}\in R^{T}$ denotes a single sample, with a label $Y^{i}\in\{1,\dots,C\}$. In Figure 1, the output shape is of dimension $C$. To formulate this better, we can write that, $p=KAN(X^{i})$ where $p=[p_{1},p_{2},\dots,p_{C}]$, is a vector of class representations.
@@
each $\Phi_{l}$ consists of a set of learnable uni-variate functions $\phi$ applied element-wise. In contrast to traditional MLPs, where nonlinearity is introduced via fixed activation functions, KANs learn the activation functions during training. A traditional MLP network performs forward computation using linear weight matrices and fixed non-linear activations:
@@
here $W_{l}$ are weight matrices and $\sigma$ denotes a fixed activation function that introduces nonlinearity into the output of each layer. KANs learn the activation functions instead of using pre-defined ones. KANs use a hybrid activation function $\phi(X)$, which combines the SiLU and spline-based basis functions. The function is defined as follows:
@@
The SiLU activation function is a continuous function, that guarantees the flow of gradient during the training process. This differentiable function has favorable optimization properties. The B-splines enable fine-grained modeling of nonlinearities through a learned basis function. Each spline function operates within a grid of size $G$:
@@
here, $B_{i}(X)$ denotes a spline basis function, $c_{i}$ represent learned coefficient or control point, $G$ is the number of grid points, and $K$ denotes the spline order (the default is $k=3$ for cubic splines). The spline order $K$ determines the smoothness of the curve and higher grid size $G$ increases the resolution of the splines. As can be seen in Equation 6, a spline with order $K$ and grid size $G$ requires $G+K$ basis functions.
@@
where, $d_{in}$ and $d_{out}$ are the input and output dimensions, $G$ is the grid size or the number of grid points, and $K$ is the spline order. Our implementation uses two stacked KAN layers. The first layer acts as the feature extractor (FE) and the second one performs the final classification. Unlike traditional MLPs, where parameter count only depends on dimensions of input and output, in KAN two other parameters, spline order and grid size, affect the number of parameters, which results in more parameters even for equivalent dimensional configurations.
@@ 4.2 Pre-processing
The only pre-processing performed in this study is just standardization with the mean ($\mu$) and standard deviation ($\sigma$) of the data. KAN has a unique ability in performing the classification of the intrusion data without the need for any sort of extra data pre-processing, which results in faster inference time. The standardization is performed as follows:
@@
we have tried min-max scaling as well, but standardization performed better than min-max scaling.
@@ 4.3 Contrastive learning
The contrastive learning framework was first introduced in [10], where it was applied in a self-supervised manner for visual representation learning. One of the key advantages of this method is that it can utilize large amount of unlabeled data created. Unlabeled data is common in various domains. This is a huge advantage, since annotating requires significant time, domain expertise, and financial resources. Contrastive learning provides a scalable solution for utilization of large amounts of unlabeled data. It pre-trains a FE, without the need for labels, and learns the underlying features present in the data. Once the pre-training is over, the parameters of the FE are frozen, and a final classifier is stacked to the network. The fine-tuning of the last layer is performed with just a fraction of the data that contains labels, without changing the parameters of the FE.
The process begins with the data $D=\{X_{scaled},Y\}$ where $X_{scaled}\in\mathbb{R}^{d_{T}}$ is standardized data, and also labels $Y$ are assumed to be unknown. Contrastive learning aims to learn similar representations for different views of the same data point that has gone through different data augmentations.
Since the similarity learning in the contrastive training is based on intra-batch distances, the first step involves sampling batches of size $M$ form the data $X_{scaled}$. Following the procedure, will result in $B(X_{scaled})$. The notation $B(\cdot)$ denotes a batch.
Now, we define the feature extractor denoted by $KAN_{\theta}(\cdot)$, with learnable parameters $\theta$, which takes an input data point $X$ and creates the representation $z$. As mentioned earlier that contrastive learning requires different views of the same data point, we propose a masking-based data augmentation strategy to randomly set features in the data point $X$ to zero. This masking function is defined as :
$x_{masked}=\text{Masking}(\cdot,p)$ (9)
where $p$ is a pre-defined masking percentage and it is set to $0.1$ in our case. The masking function randomly sets $p\%$ of the input features to zero. Following Equation 9 for an input batch twice, will result in two different batches of the same input $B(x_{scaled})$, which are, $B_{1}(X)$ and $B_{2}(X)$; note that for simplicity we remove the scaled keyword, that denoted the standard-scaled data. As mentioned before, the feature extractor $KAN_{\theta}(\cdot)$ must generate similar representations for corresponding elements of these two augmented batches. Let $B_{1}(z)$ and $B_{2}(z)$ represent the resulting batches generated from the two views, which $z$ denotes the representations created by the network, as follows:
@@
where the network $KAN_{\theta}(\cdot)$, creates a batch of representations for each input batch. The input and output batches are of size $M$. The aim is to bring the representations created for corresponding elements in the augmented data closer to each other, while distancing the dissimilar representations. First, the two batches of data must be concatenate into a single batch with $2M$ samples, as follows:
$B_{contrastive}(z)=concat(B_{1}(z);B_{2}(z))$ (12)
@@
$l_{i,j}=\log\frac{\exp(sim_{cos\_sim}(x^{i},x^{j})/\tau)}{\sum_{k=1}^{2M}1_{k\neq i}\exp(sim_{cos\_sim}(x^{i},x^{k})/\tau)}$ (13)
where $i,j\in\{1,\dots,2M\}$, denote the indices of the samples in batch $B_{contrastive}(z))$ and $sim_{cos\_sim}$ is the cosine similarity between two vectors as follows:
$sim_{cos\_sim}=\cos(\phi)=\frac{\mathbf{A}\cdot\mathbf{B}}{\|\mathbf{A}\|\|\mathbf{B}\|}$ (14)
where $\phi$ is the angle between the two latent representations. In Equation 13, the self-similarity ($i=k$) is ignored (with $1_{k\neq i}$). And $\tau$ is the temperature or scaling factor, that is set to $0.5$. $x^{i}=B_{contrastive}(z)(i)$ represents the $i^{th}$ element of the batch $B_{contrastive}(x)$, with similar definitions for $x^{j}$ and $x^{k}$.
$l$ is a $2M\times 2M$ matrix, where the elements $l(2k,2k-1)$ and $l(2k-1,2k)$ represent the pairwise similarity between the two batches. The total loss is then calculated as:
@@
Since, the denominator of Equation 13 includes all the other dissimilar samples, it captures the dissimilarity associated with the negative samples. Maximizing the distance between the dissimilar samples, will result in minimization of the loss. The parameters $\theta$ of the feature extractor are updated according to gradient descent optimization, as follows:
@@
where $\eta$ is the pre-training learning rate, $n$ denotes the $n^{th}$ iteration of training. Then the weights are saved for the fine-tuning stage. Algorithm 1 and Figure 2, show an algorithmic and a visual view of the contrastive training.
@@
3:     Create a batch of training data, with size $M$: $B(X_{scaled})$
4:     Create different views of the data with augmentations:$B_{1}(X)\leftarrow\text{Masking}(B(X_{\text{scaled}}),p)$$B_{2}(X)\leftarrow\text{Masking}(B(X_{\text{scaled}}),p)$
@@
6:     Concatenate the two batches of representations vertically to create a batch of $2M$ samples:$B_{contrastive}(z)\leftarrow concat(B_{1}(z);B_{2}(z))$
7:     Calculate the $l_{i,j}$ with the contrastive loss, and $B_{contrastive}(z)$
8:     Compute similarity loss by adding the elements of the $(2k-1,2k)$ and $(2k,2k-1)$:$Loss\leftarrow\frac{1}{2M}\sum_{k=1}^{M}(l_{2k-1,2k}+l_{2k,2k-1})$
@@
10:     Save model $KAN_{\theta}$
11:     Use $KAN_{\theta}$ to evaluate new data and use for fine-tuning

### v2

@@ 3.1 UNSW-NB15 Dataset
The UNSW-NB15 dataset, created by researchers at the University of New South Wales (UNSW), represents modern network traffic scenarios and includes a wide range of low-footprint attacks. A synthetic testbed was configured using the IXIA PerfectStorm tool to generate realistic network traffic and diverse attack scenarios. This dataset contains network traffic collected from multiple sources, including benign communications, malicious attacks, and routine operations such as software updates. Raw network packets were captured using the tcpdump tool and stored in the PCAP file format.
A total of 49 network-based features were extracted across five categories: basic features, content features, time features, connection features, and additional features such as protocol type, duration, packet/byte counts, payload analysis, jitter, inter-packet arrival times, and historical behavior. This dataset includes multiple modern attack families, including Fuzzers, Analysis, Backdoors, DoS, Exploits, Reconnaissance, Shellcode, Generic, and Worms, in addition to Normal traffic. In this work, we utilize only a subset of categories: Normal, Fuzzers, DoS, Exploits, Generic, and Reconnaissance. The dataset contains approximately 2.5 million records, consisting of 2,218,761 normal and 321,283 attack records [36].
The UNSW-NB15 dataset was created to address the limitations of previous datasets (e.g., KDD99 and NSL-KDD), such as the lack of modern low-footprint attacks and outdated normal traffic patterns. Skewness and kurtosis tests on this dataset confirm non-linearity, non-normal distributions, and feature correlations that challenge traditional machine learning methods [38]. Older datasets like KDD99 and NSL-KDD suffered from redundant records, which UNSW-NB15 aims to overcome [39, 40]. The various classes and the corresponding number of instances for each category are provided in Table 1.
It is worth noting that the original dataset contains 49 columns, including several categorical features: proto, service, and state. We apply one-hot encoding to these features, resulting in a total of 204 columns after encoding. In this work, 558,954 samples are used for the pre-training phase and 55,895 samples for the fine-tuning stage, corresponding to 22.01% and 2.20% of the total number of samples, respectively. The fine-tuning set is sampled from the pre-training set.
@@ 3.2 BoT-IoT Dataset
The BoT-IoT dataset [37], publicly released in 2018 by the Network Security Lab at the University of New Brunswick, Canada, was collected using a smart home testbed that included various IoT devices such as routers, cameras, and smart appliances. The setup was designed to mimic real-world IoT environments and includes multiple types of cyber-attacks. Network traffic was recorded over several weeks using a passive monitoring tool called CICFlowMeter, which captured data at the network gateway.
This dataset contains network features such as protocol type, source and destination IP addresses, and port numbers, with each data point labeled. It serves as a benchmark for evaluating intrusion detection systems, particularly those focused on IoT security, and includes common IoT-related attacks. The categories present in this dataset are DDoS, DoS, Normal, Theft, and Reconnaissance.
In practice, when preparing the data for training, some columns such as attack and subcategory are removed, while the category column is retained to serve as the target label. The proto column, being categorical, is transformed using one-hot encoding, similar to the preprocessing steps applied to the UNSW-NB15 dataset. After encoding, the dataset contains 20 features for training. For this dataset, we use 469,570 samples, approximately 12.8% of the total data, for the pre-training phase, while 46,957 samples, accounting for 1.28% of the total 3,668,522 samples, are employed for the fine-tuning phase. The fine-tuning subset is sampled from the pre-training subset, and labels are used only for the fine-tuning subset. Both sample sizes are significantly lower than the volumes typically required by other algorithms. Table 2 presents the classes in the BoT-IoT dataset along with the number of instances for each class.
@@ 3.3 Gas Pipeline Dataset
The Gas Pipeline dataset is developed to study intrusion detection in critical infrastructure systems that rely on SCADA technology. Its primary purpose is to simulate cyber-attacks targeting SCADA systems, serving as a standardized and common benchmark for evaluating different intrusion detection methods. This dataset is based on a small-scale gas pipeline testbed, where MODBUS traffic was recorded using a network data logger. The testbed consists of an airtight pipeline connected to a compressor, a pressure sensor, and a solenoid-controlled relief valve. The system regulates air pressure within the pipeline through a proportional-integral-derivative (PID) control scheme. In this research, we utilize the 2014 version of the dataset [35].
The dataset is composed of 27 columns, where the result column denotes the condition of the system. Therefore, the dataset contains 26 actual features related to network traffic, process control, and process measurements. Among these, seven features contain only a single constant value and do not contribute useful information to the algorithm. These features are rate, comm_write_fun, gain, reset, deadband, cycletime, and crc_rate. After dropping these features, 19 features remain for training the network. Regarding the labels, there are eight distinct conditions in the dataset: normal (Normal), Naïve Malicious Response Injection (NMRI), Complex Malicious Response Injection (CMRI), Malicious State Command Injection (MSCI), Malicious Parameter Command Injection (MPCI), Malicious Function Code Injection (MDCI), Denial of Service (DoS), and Reconnaissance (Recon). These are denoted by labels 0 to 7, respectively [35]. The classes along with their respective instance counts are shown in Table 3.
A total of 77,615 samples, representing 80% of the entire dataset, are used for the pre-training phase. From this set, 7,761 samples (8% of the total data) are selected for the fine-tuning phase. The pre-training stage does not require labels, which are only utilized during fine-tuning. This approach makes the method suitable for scenarios where abundant unlabeled data is available.
@@ 4 Methodology
In this section, we present the methodology proposed in this study. Section subsection 4.1 begins with an overview of the Kolmogorov-Arnold representation theorem and the KAN model. Then, subsection 4.2 outlines the pre-processing steps associated with the method. Finally, subsection 4.3 provides a detailed discussion of the contrastive learning approach.
@@ 4.1 KAN: Kolmogorov-Arnold Network
KAN, first introduced in [11], is an alternative to traditional fully-connected networks. It exhibits strong capabilities in modeling non-linear functions and outperforms conventional MLP layers. KAN has demonstrated significant potential in a variety of domains, including anomaly detection [41], human activity recognition [42], solving differential equations [43], and time series analysis [44]. KAN has shown notable results such as advancements in accuracy, interpretability, and computational efficiency [11].
Before proceeding to the proposed methodology, we examine the current applications of KANs in the field of anomaly detection. [45] employs KAN for CPS attack detection using a minimal and fast approach that only utilizes standard scaling; this method achieves high performance with low inference time.
In [46], the authors propose faster-KAN with the help of Gaussian radial basis functions (GRBF) to detect anomalies in IoT networks. This research is suitable for low-end devices and systems with limited computational resources. [41] introduces CCTAK, an unsupervised anomaly detection framework for IIoT networks, combining CNN and KAN. The enhanced interpretability of this method is a key advantage. [47] uses KAN for credit card fraud detection, where KAN outperforms conventional MLP networks.
In [48], the detection of botnet attacks is investigated. It is shown that the original KAN outperforms traditional ML and DL methods while also providing improved interpretability. [49] utilizes KAN for motor condition monitoring, improving real-time diagnosis of faults in motors and achieving better performance and interpretability compared to MLP networks. KAN is also adopted in the healthcare domain. In [50], the authors performed ECG-based cardiac abnormality detection, showcasing the effectiveness and generalizability of KAN for wearable medical devices.
KAN is grounded in the Kolmogorov-Arnold representation theorem, which initially served to solve differential equations. However, in this study, KAN is applied for classification tasks. The theorem guarantees that any multivariate continuous function $f(X)$, where $X=(X_{1},\dots,X_{T})$, can be decomposed into a sum of nested univariate functions. Meaning, for the smooth differentiable function $f:[0,1]^{T}\rightarrow\mathbb{R}^{d_{out}}$, there exist continuous univariate functions $\Phi_{q}$ and $\phi_{q,p}$, that satisfy the following equation:
@@
where $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\Phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. In this formulation, each $\phi_{q,p}$ and $\Phi_{q}$ is a univariate function, and the nested sum of them enables the approximation of any continuous multivariate function. The Kolmogorov-Arnold representation theorem ensures that $(2T+1)\times T$ univariate functions are enough to approximate any given $T$-variate function. Compared to Taylor or Fourier series approximations, which require an infinite number of terms to represent a multivariate function with high precision, KANs offer a significant advantage by approximating a function using a finite composition of learnable univariate functions.
KANs leverage this theoretical foundation by replacing typical pre-defined activation functions in neural networks with learned univariate functions applied to each input dimension individually. Unlike in a traditional MLP layer, where activation functions are applied to the outputs of nodes, KAN places activation functions on the edges of the network, i.e., between connected nodes. Figure 1 visualizes a KAN with an input size of $T$ and an output size of $C$. From this point on, to model the approximated function $f(\cdot)$, we use the notation $KAN(\cdot)$. Here, $X^{i}\in\mathbb{R}^{T}$ denotes a single sample, with a label $Y^{i}\in\{1,\dots,C\}$. In Figure 1, the output shape is of dimension $C$. To formulate this more clearly, we can write $p=KAN(X^{i})$ where $p=[p_{1},p_{2},\dots,p_{C}]$ is a vector of class representations.
@@
each $\Phi_{l}$ consists of a set of learnable univariate functions $\phi$ applied element-wise. In contrast to traditional MLPs, where nonlinearity is introduced via fixed activation functions, KANs learn the activation functions during training. A traditional MLP network performs forward computation using linear weight matrices and fixed nonlinear activations:
@@
here, $W_{l}$ are weight matrices and $\sigma$ denotes a fixed activation function that introduces nonlinearity into the output of each layer. KANs learn the activation functions instead of using pre-defined ones. KANs use a hybrid activation function $\phi(X)$, which combines the SiLU and spline-based basis functions. The function is defined as follows:
@@
The SiLU activation function is a continuous function that guarantees the flow of gradients during the training process. This differentiable function has favorable optimization properties. The B-splines enable fine-grained modeling of nonlinearities through a learned basis function. Each spline function operates within a grid of size $G$:
@@
here, $B_{i}(X)$ denotes a spline basis function, $c_{i}$ represents a learned coefficient or control point, $G$ is the number of grid points, and $K$ denotes the spline order (the default is $K=3$ for cubic splines). The spline order $K$ determines the smoothness of the curve, and a higher grid size $G$ increases the resolution of the splines. As shown in Equation 6, a spline with order $K$ and grid size $G$ requires $G+K$ basis functions.
@@
here, $d_{in}$ and $d_{out}$ are the input and output dimensions, $G$ is the grid size or the number of grid points, and $K$ is the spline order. Our implementation uses two stacked KAN layers. The first layer acts as the feature extractor , and the second performs the final classification. Unlike traditional MLPs, where the parameter count depends only on the input and output dimensions, in KAN two additional parameters, the spline order and the grid size, affect the number of parameters. This results in more parameters even for equivalent dimensional configurations.
@@ 4.2 Pre-processing
The only pre-processing performed in this study is standardization using the mean ($\mu$) and standard deviation ($\sigma$) of the data. KAN has a unique ability to perform classification of intrusion data without requiring any additional pre-processing, resulting in faster inference times. The standardization is performed as follows:
@@
We also experimented with min-max scaling; however, standardization outperformed min-max scaling in our experiments.
@@ 4.3 Contrastive learning
The contrastive learning framework was first introduced in [10], where it was applied in a self-supervised manner for visual representation learning. One of the key advantages of this method is its ability to utilize large amounts of unlabeled data. Such unlabeled data is common across various domains, which is a significant benefit since annotating data requires substantial time, domain expertise, and financial resources. Contrastive learning offers a scalable solution for leveraging large volumes of unlabeled data by pre-training a feature extractor without the need for labels, thereby learning the underlying features present in the data. After pre-training, the parameters of the feature extractor are frozen, and a final classifier is appended to the network. Fine-tuning is then performed only on this last layer using a small fraction of labeled data, without modifying the parameters of the FE.
The process begins with the data $D=\{X_{\text{scaled}},Y\}$ where $X_{\text{scaled}}\in\mathbb{R}^{d_{T}}$ is standardized data, and also labels $Y$ are assumed to be unknown. Contrastive learning aims to learn similar representations for different views of the same data point that has gone through different data augmentations.
Since similarity learning in contrastive training is based on intra-batch distances, the first step involves sampling batches of size $M$ from the data $X_{\text{scaled}}$. Following this procedure results in $B(X_{\text{scaled}})$, where the notation $B(\cdot)$ denotes a batch.
Now, we define the feature extractor, denoted by $KAN_{\theta}(\cdot)$ with learnable parameters $\theta$, which takes an input data point $X$ and generates its representation $z$. As mentioned earlier, contrastive learning requires different views of the same data point. To achieve this, we propose a masking-based data augmentation strategy that randomly sets some features in the data point $X$ to zero. This masking function is defined as:
$X_{\text{masked}}=\text{Masking}(\cdot,p)$ (9)
where $p$ is a pre-defined masking percentage, which is set to 0.10.10.1 in our case. The masking function randomly sets $p\%$ of the input features to zero. By applying the masking function described in Equation 9 to an input batch twice, we obtain two different augmented versions of the same input batch $B(X_{\text{scaled}})$, denoted by $B_{1}(X)$ and $B_{2}(X)$. For simplicity, we omit the scaled subscript, which previously indicated that the data was standardized.
As mentioned earlier, the feature extractor $KAN_{\theta}(\cdot)$ should produce similar representations for corresponding elements in these two augmented batches. Let $B_{1}(z)$ and $B_{2}(z)$ represent the resulting batches of learned representations obtained from each view, where $z$ denotes the output of the feature extractor. This process can be expressed as follows:
@@
where the network $KAN_{\theta}(\cdot)$ generates a batch of representations for each input batch. Both the input and output batches have a size of $M$. The objective is to bring the representations corresponding to the same original data point (across different augmentations) closer together in the representation space, while pushing apart the representations of dissimilar (non-matching) data points.
To achieve this, the two augmented batches are concatenated to form a single batch of size $2M$, as follows:
$B_{\text{contrastive}}(z)=concat(B_{1}(z);B_{2}(z))$ (12)
@@
$l_{i,j}=\log\frac{\exp(\text{sim}_{\text{cos\_sim}}(x^{i},x^{j})/\tau)}{\sum_{k=1}^{2M}1_{k\neq i}\exp(\text{sim}_{\text{cos\_sim}}(x^{i},x^{k})/\tau)}$ (13)
where $i,j\in\{1,\dots,2M\}$, denote the indices of the samples in batch $B_{\text{contrastive}}(z))$ and $\text{sim}_{\text{cos\_sim}}$ is the cosine similarity between two vectors as follows:
$\text{sim}_{\text{cos\_sim}}=\cos(\phi)=\frac{\mathbf{A}\cdot\mathbf{B}}{\|\mathbf{A}\|\|\mathbf{B}\|}$ (14)
where $\phi$ is the angle between the two latent representations. In Equation 13, the self-similarity ($i=k$) is ignored using the indicator function $1_{k\neq i}$. The term $\tau$ is the temperature (or scaling) factor, which is set to 0.50.50.5. The notation $x^{i}=B_{\text{contrastive}}(z)(i)$ represents the $i^{\text{th}}$ element of the batch $B_{\text{contrastive}}(z)$, with similar definitions for $x^{j}$ and $x^{k}$.
$l$ is a $2M\times 2M$ matrix, where the elements $l(2k,2k-1)$ and $l(2k-1,2k)$ represent the pairwise similarities between the two augmented batches. The total loss is then calculated as:
@@
Since the denominator of Equation 13 includes all the dissimilar samples, it captures the dissimilarity associated with the negative samples. Maximizing the distance between dissimilar samples results in minimizing the loss. The parameters $\theta$ of the feature extractor are updated using gradient descent optimization, as follows:
@@
where $\eta$ is the pre-training learning rate, and $n$ denotes the $n^{th}$ iteration of training. The weights are then saved for the fine-tuning stage. Algorithm 1 and Figure Figure 2 present an algorithmic and visual overview of the contrastive training process, respectively.
@@
3:  Create a batch of training data, with size $M$: $B(X_{\text{scaled}})$
[…]
@@
[…]
7:  Calculate the $l_{i,j}$ with the contrastive loss, and $B_{\text{contrastive}}(z)$
[…]
@@
10:  Save model $KAN_{\theta}(\cdot)$
11:  Use $KAN_{\theta}(\cdot)$ to evaluate new data and use for fine-tuning